# Quantitative Evaluation of Multiple Instance Learning Reliability For WSIs Classification

### v2

@@ 3.5 Conclusion
3.6 Acknowledgments
This work was done as a part of the IMI BigPicture project (IMI945358).
Model Reliability Classification Computation MI Spearman’s AUPRC AUC F1 FLOPs Model Size ABMIL-ADD-ATT 0.2063 0.4178 0.7770 0.918 0.840 94.6 M 789.5 K ABMIL-ADD-PATCH 0.0571 0.3035 0.7045 CLAM-ADD-ATT 0.1884 0.4324 0.7789 0.921 0.834 94.6 M 793.3 K CLAM-ADD-PATCH 0.0539 0.2908 0.6992 DTFD-PATCH 0.0651 0.3343 0.7309 0.947 0.882 125.9 M 1315.7 K ACMIL-ADD-ATT 0.1755 0.3724 0.7537 0.905 0.856 94.6 M 796.5 K ACMIL-ADD-PATCH 0.0858 0.1739 0.6227 MAX-POOL-INS 0.0003 0.0206 0.4921 0.882 0.791 63.1 M 526.6 K MAX-POOL-INS-PATCH 0.2262 0.4049 0.7564 MEAN-POOL-INS 0.2987 0.4298 0.8079 0.846 0.749 63.1 M 526.6 K

### v3

@@ Conclusion
Supporting information
S1 Fig. Learning curve of the models (set 1). Each panel shows the training curve for one model on TCGA BRCA.
S2 Fig. Learning curve of the models (set 2). Each panel shows the training curve for one model on TCGA BRCA.
Data Availability Statement
The code for reproducing our results is available at https://github.com/tueimage/MIL-Reliability.
Competing interests
The authors have declared that no competing interests exist. This does not alter our adherence to PLOS ONE policies on sharing data and materials.
Author Contributions
Conceptualization: Hassan Keshvarikhojasteh, Mitko Veta
Data curation: Marc Aubreville, Christof A. Bertram
Formal analysis: Hassan Keshvarikhojasteh
Methodology: Hassan Keshvarikhojasteh, Mitko Veta
Supervision: Mitko Veta, Josien P.W. Pluim
Writing – original draft: Hassan Keshvarikhojasteh
Writing – review & editing: Hassan Keshvarikhojaste, Marc Aubreville, Christof A. Bertram, Josien P.W. Pluim, Mitko Veta
Funding
This research was supported by the IMI BigPicture project (IMI945358), part of the Innovative Medicines Initiative 2 Joint Undertaking under grant agreement No 945358. The funders had no role in study design, data collection and analysis, decision to publish, or preparation of the manuscript.